Viewing Report – 9th May 2020 – IMT3 – Dark frames, filter

22:05 Frame and focus on 9.26 and 7.64 magnitude stars used before moving to M61 to capture some Lum frames for calibration of flats to solve the doughnut embossing.

22:18 Slewed to M61, performed Solve ‘n’ sync, slew here for centring the object, ran four Luminance subs of 300 seconds each (Bin1x1) at Gain 139 and Offset 21. This was completed by 22:38.

22:45 Chief TOSA then warmed up the CMOS camera, set the filter wheel to the empty slot position, powered off and disconnected the camera and filter wheel. This was so the ASI1600MM and filter wheel could be detached and a blanking filter installed into the empty filter wheel slot position without having to open up the filter wheel. The idea is that we move to the blank filter position when taking Dark subs to prevent light leakage on the Officina Stellare 12-inch.

22:55 Everything was back online and the SGPro profiles were modified to reflect the new blank filter location.

23:13 Unfortunately a dust mote was introduced onto the CMOS sensor window so we sent Chief TOSA back out to the dome to do the job properly this time 🙂 This meant parking the mount, warming up the camera, powering off the camera, remove it, clean the sensor and reattach the camera.

23:24 All reassembled and powered back on, slewed to NGC 4147 (Globular cluster) ready for a a 60 second Lum filter test which showed that a decent cleaning job had been done ….. about time too !

23:40 As it was getting hazy/cloudy it was decided to collect some calibration frames, in this case 25×10 minute dark subs at Gain 39 Offset 21 using the new blank filter.

23:50 We logged off from the remote session and left the dark frame sequence running until it was due to finish in the early morning.

Almost Completed

So another day spent with Dave and Bob just finishing off the finishing touches to the scope/mount installation whilst Dave messed about pretending to be a brickie and spent lots of time not actually doing any brick work so we think he may now be qualified.

We spent some time tidying up cabling and testing a low light camera for use in the dome at night so we can see what the mount and dome are actually doing.

Now the Chroma Ha and OIII 3nm filters are now installed into the SX filter wheel the only things that remains is to wait for the delivery of a cable (FTDI) to connect to the Optec Flip Flat panel.

Obviously the next bit is working out if the reducer to CCD spacing for the 3rd scope is correct and work out the focuser backlash so we can correctly use SGPro focus offsets for the filters.